40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3391     Entry time: Tue Aug 10 05:56:07 2010
Author: Koji 
Type: Update 
Category: PSL 
Subject: PSL Table Lifting Preparation 

Work on Aug 9th

Steve, Jenne, Koji, Alberto, Aidan, Jan, Sharmila, Katherine

From 9am to 6pm

  • Shutting down the PSL after the 90885 hours of service
  • Removals
    • The accelerometers
    • The reference cavity chamber
    • The periscopes were left so far
  • The cables between the PSL table and the outside have been disconnected
    • Listed items on Friday
    • Some unlisted items recorded and disconnected
    • Drained the cooling water from the chiller lines
    • Pulled out the chiller connections at the control room as well as the chiller control cables (temp sens & RS232C)
  • Sealed the PSL table with plastic sheets
    • Put antistatic films to the table (they are supported by the long optical poles)
    • Used capton tapes to fix the films on to the table
    • Put the white huge plastic sheet to cover the table at once
    • Some spaces at the edge of the tables are left flat such that the C-clamps can be attached
  • Sealed the AP and the ITMY tables by capton tapes

Some photos are attached in this entry. All of the photos found in the picasa album (click the slideshow)

Attachment 1: IMG_2686.jpg  207 kB  Uploaded Tue Aug 10 08:01:34 2010  | Hide | Hide all
IMG_2686.jpg
Attachment 2: IMG_2691.jpg  227 kB  Uploaded Tue Aug 10 08:01:53 2010  | Hide | Hide all
IMG_2691.jpg
Attachment 3: IMG_2692.jpg  171 kB  Uploaded Tue Aug 10 08:02:08 2010  | Hide | Hide all
IMG_2692.jpg
Attachment 4: IMG_2711.jpg  109 kB  Uploaded Tue Aug 10 08:02:25 2010  | Hide | Hide all
IMG_2711.jpg
ELOG V3.1.3-